OpenStudy (anonymous):

will fan and give a medal hurry quick come and answer this easy question if a train travels at distance of 2.50 miles in 1/3 hour of an hour . how far can the train travel in 3 hours

OpenStudy (anonymous):

22.5

OpenStudy (anonymous):

can the train travel in 3 hours=22.5

OpenStudy (anonymous):

how u know

OpenStudy (anonymous):

you know/: 2.5 miles 1/3 h X 3h then we have find the x=2.5 * 3 / (1/3) = 22.5

OpenStudy (anonymous):

then how many miles can the train go per hour

OpenStudy (anonymous):

shouldn't it be 2.50 times 3 since 3 is the recipricol of 1/3

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

k

OpenStudy (anonymous):

the ans is true:

OpenStudy (anonymous):

what that's not the q was

OpenStudy (anonymous):

the questions was it was how many miles per hour can the train travel

OpenStudy (anonymous):

oh r u talking about the 3 miles one

OpenStudy (anonymous):

ok, no this is a unit it means miles/hour

OpenStudy (anonymous):

yas

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

so now it would be 2.50 times 1/3

OpenStudy (anonymous):

no 2.5 times 3

OpenStudy (anonymous):

which is 7.5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is taht right

OpenStudy (anonymous):

just devide in (1/3) then you have: x=22.5

OpenStudy (anonymous):

waith whats the ans to #1 i no the ans to #2 is 7.5

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

the ans to #1

OpenStudy (anonymous):

heeelloooo??

OpenStudy (anonymous):

...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hurry up

OpenStudy (anonymous):

the ans is true. what exactly do you want?

OpenStudy (anonymous):

the ans to this how many miles can the train travel in 3 hours

OpenStudy (anonymous):

ok, the ans is 22.5 miles per hour train can travel. it's easy

OpenStudy (anonymous):

alright sir u sure